# IFT-UAM/CSIC-24-92

### v1

@@ Abstract
We study how tiny Dirac neutrino masses consistent with experimental constraints can arise in string theory SM-like vacua. We use as a laboratory 4d ${\cal N}=1$ type IIA Calabi–Yau orientifold compactifications, and in particular recent results on Yukawa couplings at infinite field-space distance. In this regime we find Dirac neutrino masses of the form $m_{\nu}\simeq g_{\nu}\langle H\rangle$, with $g_{\nu}$ the gauge coupling of the massive $U(1)$ under which the right-handed neutrinos $\nu_{R}$ are charged, and which should be in the range $g_{\nu}\simeq 10^{-14}-10^{-12}$ to reproduce neutrino data. The neutrino mass suppression occurs because the right-handed neutrino kinetic term behaves as $K_{\nu\nu}\simeq 1/g_{\nu}^{2}$. At the same time a tower of $\nu_{R}$-like states appears with characteristic scale $m_{0}\simeq g_{\nu}^{2}M_{\rm P}\simeq 0.1-500$ eV, in agreement with Swampland expectations. Two large hidden dimensions only felt by the $\nu_{R}$ sector arise at the same scale, while the string scale is around $M_{s}\simeq g_{\nu}M_{\rm P}\simeq 10-700$ TeV. Some phenomenological implications and model building challenges are described. As a byproduct, independently of the neutrino issue, we argue that a single large dimension in the context of SM-like type IIA Calabi–Yau orientifolds leads to too small Yukawa couplings for quarks and charged leptons.
@@ 1 Introduction
$Y_{ijk}\,=\,\frac{W_{ijk}}{{\rm Vol}_{X}^{1/2}}e^{\phi_{4}/2}\Theta_{ijk}^{1/4}\,.$ (1.4)
On the one hand, ${\rm Vol}_{X}$ is the volume of the compact manifold in units of $M_{s}$ and $W_{ijk}$ is the holomorphic superpotential, both depending only on the complexified Kähler moduli $T^{a}$. On the other hand $\phi_{4}$ is the 4d dilaton and the functions $\Theta_{ijk}$ encode the information about the intersection angles of the three D6-branes involved in the Yukawa coupling, both quantities depending only on the complex structure fields $u^{K}$. As pointed out in [17], the interesting limit for us will be the one in which we keep the Kähler moduli fixed and ${\rm Vol}_{X}$ is not too large, since the ${\rm Vol}_{X}\rightarrow\infty$ corresponds to an M-theory limit in which we are not interested. Alternatively, perturbative limits with vanishing Yukawas imply some complex structure fields becoming large. In a limit parametrised by a single complex structure direction $u\rightarrow\infty$, Yukawa couplings behave like $Y\simeq 1/u^{r}$ with $r$ some rational number typically in the range $1/4\leq r\leq 1$ [17]. At the same time towers of light states with the same quantum numbers as the massless quarks and leptons, termed ‘gonions’ in [18], become light.
@@
The results obtained for the mass scales and towers of states in these types of compactifications are, as expected, consistent with general Swampland arguments (see [21, 22, 23, 24] for reviews). In particular light towers of states are obtained along large complex structure limits, in agreement with the Swampland Distance Conjecture (SDC), and such towers obey the Weak Gravity Conjecture (WGC) for the $U(1)_{\nu}$ gauge boson coupling to $\nu_{R}$’s. In fact there are arguments, based on the AdS instability conjecture [25, 26] which states that neutrinos must be Dirac and the lightest neutrino must have a mass bounded by the cosmological constant scale, $m_{\nu}^{\rm min}\lesssim\Lambda_{{\rm cc}}^{1/2}$ [27, 28, 29]. This nicely fits with the Dirac option which we analyse in this paper. We describe some consequences of these Swampland arguments in section 5. In combination with our previous results, they give a rationale for why two large dimensions should open: they are required to obtain a sufficiently light Dirac neutrino to obey the Swampland bound $m_{\nu}^{\rm min}\lesssim\Lambda_{{\rm cc}}^{1/4}$. This also explains the apparent coincidence between the neutrino mass scale and that of the cosmological constant.
@@
Some technical details have been relegated to three appendices. In appendix A we explain the difficulties of having a SM-like orientifold with a single large dimension and only the neutrino Yukawa couplings small, and then rule out settings with more than two large dimensions because they imply a very small string scale. In appendix B we present an $SU(5)$ intersection brane setting in which the same general structure for obtaining tiny Yukawa couplings is realised. Appendix C contains an analysis of the $U(1)$ anomalies and the conditions for $U(1)$’s other than hypercharge becoming massive in the explicit SM-like configuration described in the main text.
@@ 2.2 Yukawas at infinite distance
The challenge that we address in this work is to identify moduli space directions in which one can suppress the Yukawa couplings of neutrinos while maintaining unsuppressed (or not exceedingly suppressed) the rest of the SM Yukawa couplings. In this sense, notice that taking the limit ${\rm Vol}_{X}\rightarrow\infty$ is not a good strategy, since it affects all the Yukawas in a universal way. Furthermore, if we take this limit at constant or slowly varying $e^{\phi_{4}}$, then a KK-like tower of D0-branes becomes light, signalling a decompactification regime in which one extra dimension opens up, and the microscopic physics of the system is better described by M-theory on a $G_{2}$-manifold. [^4] In type IIA CY orientifolds, the D0-brane tower is made of non-BPS unstable states, which are however very long-lived when located at a point $p\in X$ far away from the O6-plane. From here one expects that at large $g_{s}$, $M_{s}/g_{s}$ controls the scale of a KK tower in the M-theory $G_{2}$ manifold. A different strategy to get almost vanishing Yukawa couplings is to construct models in which the holomorphic Yukawas $W_{ijk}$ have some zero entries, which then get corrected by exponentially suppressed contributions from Euclidean E2-brane charged instantons, see refs.[13, 14, 15, 16, 48, 49, 50, 51]. However, in practice it is quite challenging to find explicit D6-brane models in which all neutrino Yukawas become very small while avoiding that some other quarks or charged lepton Yukawas also become small, see some attempts in [48, 49, 51] and references therein.
@@ 2.2.2 Difficulties for a single large dimension or more than two
In the above $s,u^{(2)}\rightarrow\infty$ limit two dimensions become large. An interesting question is how things change if only one dimension is large. One can see that this is obtained in the limit $s,u^{(2)},t^{(2)}\sim u$, $u\rightarrow\infty$, as explained in appendix A. Then the winding states remain at the string scale and only the KK tower becomes light. This however turns out to be problematic for the following reason. All Yukawa couplings in the theory get an overall prefactor ${\rm Vol}_{X}^{-1/2}$. Thus in a theory containing the SM in the limit $t^{(2)}\rightarrow\infty$ the Yukawas of all SM particles will get suppressed, and not only those of neutrinos. We know however that there are Yukawa couplings of order one, like that of the top quark, so that this limit is inconsistent with observation.
Notice that this unwanted suppression is there irrespective of what the origin of tiny neutrino masses is. Thus a single large dimension implies in general too small Yukawa couplings. This is certainly true in the toroidal case. One may argue that in a general CY setting this need not be the case since in principle one can have that ${\rm Vol}_{X}$ remains unchanged even when a specific Kähler modulus grows to infinity. However, as we discuss in appendix A, this is only possible at the cost of some other Kähler modulus getting small. Then D2-branes wrapping a two-cycle whose area is controlled by this modulus will give rise to a light tower of particles, which will then open an additional dimension so that actually we no longer are in a single tower regime. Alternatively one could have D4-branes wrapping a shrinking four-cycle that generate a light tower with an analogous effect. Similar considerations can be drawn whenever one considers setups with an odd number of large dimensions. The reader interested in more details about the case with odd extra dimensions may consult appendix A.
Finally, one may consider limits in which four or six dimensions become large. This possibility does not imply having suppressed SM Yukawa couplings but, as discussed in appendix A, when applied to scenarios with small Dirac neutrino masses like the one in section 3, it gives rise to a quantum gravity cut-off that is ruled out experimentally. Given these difficulties, in the rest of the paper we will focus on setups with two large geometric dimensions, as in the toy model above.
@@ 6 Summary and outlook
We have also examined the case of a single large dimension within the context of type IIA CY orientifolds with SM content. We find that irrespective of the issue of neutrino masses, in such a case all Yukawa couplings (not only the neutrino ones) are suppressed, which would be inconsistent with experiment.
@@ Appendix A The case of a single large dimension and Yukawa couplings
In the field space limits considered in the main text, we only treat the case in which two dimensions of the compact manifold $X_{6}$ are decompactified simultaneously. We consider here the case in which a single dimension of $X_{6}$ becomes large and also the intermediate cases in which two dimensions decompactify at two different scales. As we will explain below, although these limits are perfectly consistent and give rise to suppressed Yukawa couplings, we will argue they are not acceptable since one cannot avoid getting too small Yukawa couplings for all SM fermions, and not only for neutrinos. Moreover, when one considers more than two dimensions either the same problem is faced or our scheme provides a string scale so low that is ruled out by current experimental data.
The orientifold toy model with only one large dimension
Let us first consider limits with one large extra dimension in the ${\bf Z}_{2}\times{\bf Z}_{2}$ toy model of section 2.2.1 with gauge group $U(4)\times SU(2)_{R}\times SU(2)_{L}$ described in the main text. Let us consider again the $s,u^{(2)}=u$, $u\rightarrow\infty$ limit, which we need in order to get small Yukawas. In this second torus, we have Kaluza–Klein and winding states
$m_{\rm KK}\ =\ \frac{2M_{s}}{R_{3}}\ =\ \frac{M_{\rm P}}{2A_{2}^{1/2}\sqrt{su^{(2)}}}\,,\qquad m_{\rm w}\ =\ \frac{1}{2}R_{4}M_{s}\ =\ \frac{A_{2}^{1/2}M_{\rm P}}{2\sqrt{su^{(2)}}}\,.$ (A.1)
@@
We now consider the limit in which $R_{3}\ \sim u\rightarrow\infty$ (implying $s,u^{(2)}\rightarrow u$, with $g_{s}=e^{\phi}$ constant), keeping $R_{4}$ constant, so that we are left with only one large dimension. Since $\mbox{Im}\,T_{2}=t^{(2)}\equiv t=A_{2}$, we need to take the limit:
$s,u^{(2)}\sim\,u\,,\quad t\,\sim\,u\,.$ (A.2)
@@
There are several quantities that only depend on the complex structure, like the gauge kinetic function, FI-terms, and the gonion masses. Thus one still has
$g^{2}\equiv g_{a}^{2}\simeq g_{R}^{2}\sim\frac{1}{u}\,,\qquad m_{\rm gon}\ \sim\frac{M_{\rm P}}{u}\,,$ (A.3)
just like the limits in the main text. However, the Yukawa couplings will change. Given that
$Y_{ijk}\ =\ \frac{W_{ijk}}{{\rm Vol}_{X}^{1/2}}e^{\phi_{4}/2}\Theta_{ijk}^{1/4}\ ,$ (A.4)
writing ${\rm Vol}_{X}={\rm Vol}_{4}\times A_{2}$, in the limit (A.2) we get
$Y_{{F_{R}}F_{L}H}\,\sim\,\frac{1}{t^{1/2}}\left(\frac{m_{{\rm gon},2}}{M_{\rm P}}\right)^{1/2}\sim\frac{1}{u}\,\sim\,g^{2}\,.$ (A.5)
@@
This is at variance with the two-large dimension case in which $Y\sim g$. Also at variance is the scale of the large dimension $R_{3}$ which in this case is given by
$m_{\rm KK}\,\sim\,\frac{M_{\rm P}}{t^{1/2}u}\,\sim\,\frac{M_{\rm P}}{u^{3/2}}\,\sim\,g^{3}M_{\rm P}\,,$ (A.6)
@@
whereas in the case of two large dimensions we have $m_{\rm KK}\sim g^{2}M_{\rm P}$. Note that in the case of two large dimensions KK, winding and gonion masses are of the same order. In the case of just one large dimension, the gonion masses are $g^{-1}$ times larger. The string scale in this case is still $M_{s}=e^{\phi_{4}}M_{\rm P}\sim M_{\rm P}/u^{1/2}$. It corresponds to the species scale $\Lambda_{\rm QG}$ that one obtains with a single KK tower:
$\Lambda_{\rm QG}\,\sim(m_{\rm KK})^{1/3}M_{\rm P}^{2/3}\,\sim\,gM_{\rm P}.$ (A.7)
All in all the structure of scales is as shown in figure 5.
Figure 5: Structure of scales for one large dimension.
Application to small neutrino masses
@@
One can translate this structure to a brane setting hosting the SM. The main point is that by imposing that the neutrino Yukawa has the required size, one fixes the value of $g_{\nu}\simeq 1/u^{1/2}$:
@@
This value for $g_{\nu}$ is much larger than in the case of two large dimensions, which led to $g_{\nu}\simeq 10^{-12}$. Then numerically the scales are like in table 4.
String scale SM gonions $\nu_{R}$ tower large dim Vector boson
$M_{s}$ $m_{\rm gon}^{\rm SM}$ $m_{\rm gon}^{\nu}$ $m_{\rm KK}$ $M_{V_{\nu}}$
$g_{\nu}M_{\rm P}$ $\lesssim g_{\nu}M_{\rm P}$ $g_{\nu}^{2}M_{\rm P}$ $g_{\nu}^{3}M_{\rm P}$ $\geq g_{\nu}^{3}M_{\rm P}$
$10^{12}$ GeV $\lesssim 10^{12}$ GeV $10^{6}$ GeV $1$ GeV $\geq 1$ GeV
Table 4: Spectrum of masses and scales in the case of one large dimension
Although this setting looks attractive, it has a lethal problem in its application to the smallness of neutrino masses. In the expression for Yukawas the suppression factor ${\rm Vol}_{X}^{-1/2}$ is what gives us the extra factor $t^{-1/2}$ in the Yukawas, which makes possible the presence of a single dimension. However that factor is universal, and should be present for any Yukawa couplings, also for the normal Yukawas that we do not want to be too small (e.g. the top Yukawa is of order one). So to avoid this unwanted suppression we cannot allow $t$ grow with $u$. So only the case with two dimensions seems to be viable at the phenomenological level.
The dark dimension scenario and tiny Yukawa couplings
In the dark dimension scenario [80, 81, 82] there is only one extra dimension and one imposes that $m_{\rm KK}\simeq V^{1/4}\simeq 10^{-12}$ GeV. The motivation is not to get an explanation for why neutrino masses are so small, but instead the Swampland conjecture which states that in the limit in which the cosmological constant goes to zero, a tower of KK-like states should become massless [83]. An interesting question is whether this kind of tower is consistent with a small neutrino Yukawa coupling as we study in the present work. Using the above formulae for a single large dimension, one finds that

### v2

@@ Abstract
We study how tiny Dirac neutrino masses consistent with experimental constraints can arise in string theory SM-like vacua. We use as a laboratory 4d ${\cal N}=1$ type IIA Calabi–Yau orientifold compactifications, and in particular recent results on Yukawa couplings at infinite field-space distance. In this regime we find Dirac neutrino masses of the form $m_{\nu}\simeq g_{\nu}\langle H\rangle$, with $g_{\nu}$ the gauge coupling of the massive $U(1)$ under which the right-handed neutrinos $\nu_{R}$ are charged, and which should be in the range $g_{\nu}\simeq 10^{-14}-10^{-12}$ to reproduce neutrino data. The neutrino mass suppression occurs because the right-handed neutrino kinetic term behaves as $K_{\nu\nu}\simeq 1/g_{\nu}^{2}$. At the same time a tower of $\nu_{R}$-like states appears with characteristic scale $m_{0}\simeq g_{\nu}^{2}M_{\rm P}\simeq 0.1-500$ eV, in agreement with Swampland expectations. Two large hidden dimensions only felt by the $\nu_{R}$ sector arise at the same scale, while the string scale is around $M_{s}\simeq g_{\nu}M_{\rm P}\simeq 10-700$ TeV. Some phenomenological implications and model building challenges are described. We also describe the difficulties in obtaining appropriate tiny neutrino Yukawas in the cases of a single or more than two large dimensions. Thus the case with two large dimensions seems to be quite unique.
@@ 1 Introduction
$Y_{ijk}\,=\,e^{\phi_{4}/2}{\rm Vol}_{X}^{1/4}W_{ijk}\Theta_{ijk}^{1/4}\,.$ (1.4)
On the one hand, ${\rm Vol}_{X}$ is the volume of the compact manifold in units of $M_{s}$ and $W_{ijk}$ is the holomorphic superpotential, both depending only on the complexified Kähler moduli $T^{a}$. On the other hand $\phi_{4}$ is the 4d dilaton and the functions $\Theta_{ijk}$ encode the information about the intersection angles of the three D6-branes involved in the Yukawa coupling, both quantities depending only on the complex structure fields $u^{K}$. As pointed out in [17], the interesting limit for us will be the one in which we keep the Kähler moduli fixed and ${\rm Vol}_{X}$ is not too large. In that case, limits with vanishing Yukawas imply some complex structure fields becoming large. In a limit parametrised by a single complex structure direction $u\rightarrow\infty$, Yukawa couplings behave like $Y\simeq 1/u^{r}$ with $r$ some rational number typically in the range $1/4\leq r\leq 1$ [17]. At the same time towers of light states with the same quantum numbers as the massless quarks and leptons, termed ‘gonions’ in [18], become light.
@@
The results obtained for the mass scales and towers of states in these types of compactifications are, as expected, consistent with general Swampland arguments (see [21, 22, 23, 24] for reviews). In particular light towers of states are obtained along large complex structure limits, in agreement with the Swampland Distance Conjecture (SDC), and such towers obey the Weak Gravity Conjecture (WGC) for the $U(1)_{\nu}$ gauge boson coupling to $\nu_{R}$’s. In fact there are arguments, based on the AdS instability conjecture [25, 26] which state that neutrinos must be Dirac and the lightest neutrino must have a mass bounded by the cosmological constant scale, $m_{\nu}^{\rm min}\lesssim\Lambda_{{\rm cc}}^{1/2}$ [27, 28, 29]. This nicely fits with the Dirac option which we analyse in this paper. We describe some consequences of these Swampland arguments in section 5. In combination with our previous results, they give a rationale for why two large dimensions should open: they are required to obtain a sufficiently light Dirac neutrino to obey the Swampland bound $m_{\nu}^{\rm min}\lesssim\Lambda_{{\rm cc}}^{1/4}$. This also explains the apparent coincidence between the neutrino mass scale and that of the cosmological constant.
@@
Some technical details have been relegated to three appendices. In appendix A we explain the difficulties of having a SM-like orientifold with a single large dimension and neutrino Yukawa couplings small, and then rule out settings with more than two large dimensions because they imply a very small string scale. In appendix B we present an $SU(5)$ intersection brane setting in which the same general structure for obtaining tiny Yukawa couplings is realised. Appendix C contains an analysis of the $U(1)$ anomalies and the conditions for $U(1)$’s other than hypercharge becoming massive in the explicit SM-like configuration described in the main text.
@@ 2.2 Yukawas at infinite distance
The challenge that we address in this work is to identify moduli space directions in which one can suppress the Yukawa couplings of neutrinos while maintaining unsuppressed (or not exceedingly suppressed) the rest of the SM Yukawas. In this sense, taking the limit $e^{\phi_{4}}{\rm Vol}_{X}^{1/2}=g_{s}\to 0$ is not a good strategy, since this factor affects all the Yukawas in a universal way. A different procedure to get almost vanishing Yukawa couplings for the neutrinos is to construct models in which, by selecting certain points in Kähler moduli space the holomorphic Yukawas $W_{ijk}$ have some zero entries, which then get corrected by exponentially suppressed contributions from Euclidean E2-brane charged instantons, see refs.[48, 49, 50, 51, 52, 53, 54]. However, in practice it is quite challenging to find explicit D6-brane models in which all neutrino Yukawas become very small while avoiding that some other quarks or charged lepton Yukawas also become small, see some attempts in [50, 51, 54] and references therein.
@@ 2.2.2 Difficulties for a single large dimension or more than two
In the above $s,u^{(2)}\rightarrow\infty$ limit two dimensions become large. An interesting question is how things change if only one dimension is large. One can see that this is obtained in the limit $s,u^{(2)},t^{(2)}\sim u$, $u\rightarrow\infty$, as explained in appendix A. Then the winding states remain at the string scale and only the KK tower becomes light. As we describe in that appendix, getting a Yukawa coupling for neutrinos of the required size, implies the extra dimension must be extremely large, with $m_{KK}\simeq 10^{-57}$ GeV and a nonsensical Species Scale of order 100 eV. So making a single dimension large is not the way to get tiny Dirac neutrino masses of the correct size.
A single large extra dimension has also received renewed interest within the context of the idea of ‘Dark Dimension’ [57, 58, 59, 60]. In this case one has $m_{\rm KK}\simeq 10^{-3}$ eV, in possible connection with the scale of the cosmological constant. As we said, as in all one large dimension models, one cannot obtain the required Dirac neutrino masses from tiny Yukawas. On the other hand we also show in Appendix A that such a large dimension is compatible with the correct overall size for the SM Yukawa couplings only if along the large dimension there are no light SM gonion towers.
Finally, one may consider limits in which four or six dimensions become large, see Appendix A. When applied to scenarios with small Dirac neutrino masses like the one in section 3, it gives rise to a too low quantum gravity cut-off that is ruled out experimentally. Given these difficulties, in the rest of the paper we will focus on setups with two large geometric dimensions, as in the toy model above.
@@ 6 Summary and outlook
We have also examined the case of a single large dimension within the context of type IIA CY orientifolds with SM content. We find that one large extra dimension cannot explain the tiny value of the Dirac neutrino. Moreover, regardless of the question of neutrino masses, it could be compatible with the existence of a realistic Yukawa coupling structure only if there is no light SM gonion tower along the large dimension. On the contrary, a light gonion tower along it would produce very suppressed Standard Model couplings. Further, we have also considered the possibility that four or six dimensions become large, and when applied to scenarios with tiny Dirac neutrino masses, result in a quantum gravity cutoff that is too low and already experimentally ruled out.
@@ Appendix A The case of a single large dimension and Yukawa couplings
We have seen that starting with a 4D $N=1$ SUSY vacuum, obtaining small neutrino Yukawa couplings naturally leads to two decompactified dimensions. It is interesting to see however whether the case with a single large dimension and a small neutrino Yukawa coupling is also viable.
Let us consider as a guideline the ${\bf Z}_{2}\times{\bf Z}_{2}$ toy model with gauge group $U(4)\times Sp(2)_{R}\times Sp(2)_{L}$ described in the main text. Now, we know that in the i-th torus we have KK and winding states, given by
$m_{{\rm KK},i}\,=\,\frac{2M_{s}}{R_{2i-1}}\,=\,\frac{M_{\rm P}}{2A_{i}^{1/2}\sqrt{su^{(i)}}}\,,\qquad m_{{\rm w},i}\,=\,\frac{1}{2}R_{2i}M_{s}\,=\,\frac{A_{i}^{1/2}M_{\rm P}}{2\sqrt{su^{(i)}}}\,.$ (A.1)
@@
We want to take the limit in which $R_{2i-1}\sim u\rightarrow\infty$ for some $i$, (implying $s,u^{(i)}\rightarrow u$, with $e^{-\phi}$ invariant) while keeping $R_{2i}$ constant in string units, so we are left with only one large dimension. Since ${\rm Im}T_{i}=t^{(i)}=t=A_{i}$, we need to take the limit:
$s,u^{(i)}\ \sim\ u\ ;\ t^{(i)}\sim\ u\ ;\ u\longrightarrow\ \infty\ .$ (A.2)
One then has for the KK and winding scales along the i-th plane the following scaling
$m_{{\rm KK},i}\ \simeq\ \frac{M_{\rm P}}{u^{3/2}}\ ;\ m_{{\rm w},i}\ \ \simeq\ M_{s}\ \simeq\ \frac{M_{\rm P}}{u^{1/2}}.$ (A.3)
In order to obtain some tiny Yukawa coupling we need to take the limit along a complex plane in which that limit implies a light tower of gonions. In the toy example that means either $i=2$ or $i=3$. So let us consider for definiteness $i=2$.
@@
There are several quantities that only depend on the complex structure. In particular the gauge kinetic function and the FI-terms. Also the gonion masses. Thus one has:
$g^{2}\equiv g_{a}^{2}\ \simeq\ g_{R}^{2}\ \simeq\frac{1}{u}\ ;\ m_{gon}\ \simeq\ \frac{M_{\rm P}}{u}\ .$ (A.4)
Recalling the expression for Yukawa couplings
$Y_{ijk}\,=\,e^{\phi_{4}/2}{\rm Vol}_{X}^{1/4}W_{ijk}\Theta_{ijk}^{1/4}\,,$ (A.5)
and writing ${\rm Vol}_{X}={\rm Vol}_{4}\times A_{2}$, in the studied limit we get
$Y_{{F_{R}}F_{L}H}\ \simeq\ t^{1/4}\left(\frac{m_{{\rm gon},2}}{M_{\rm P}}\right)^{1/2}\ \simeq\ \frac{t^{1/4}}{u^{1/2}}\ \simeq\ g^{1/2}$ (A.6)
@@
This is at variance with the 2-large dimension case in which $Y\sim g$. Also at variance is the scale of the large dimension $R_{3}$ which in this case is given by
$m_{\rm KK}\ \simeq\ \frac{M_{\rm P}}{t^{1/2}u}\ \simeq\frac{M_{\rm P}}{u^{3/2}}\ \simeq\ g^{3}M_{\rm P},$ (A.7)
@@
whereas in the case of 2 large dimensions is $m_{\rm KK}\sim g^{2}M_{\rm P}$. Note that in the case of 2 large dimensions ${\rm KK}$, winding and gonion masses are of the same order. In the case of 1 large dimension the gonions masses are $g^{-1}$ times larger. The string scale in this case is still $M_{s}=e^{\phi_{4}}M_{\rm P}\simeq M_{\rm P}/(u^{1/2})$. It corresponds to the species scale since:
$\Lambda\ \simeq(m_{\rm KK})^{1/3}M_{\rm P}^{2/3}\ \simeq gM_{\rm P}.$ (A.8)
A.1 Application to a small Yukawa neutrino setting
@@
One can translate this structure to e.g. the brane setting with the SM. Now the point is that imposing that the Yukawa has the required size one fixes the value of $g_{\nu}\simeq 1/u^{1/2}$:
@@
This value for $g_{\nu}$ is much smaller than in the case of 2 large dimensions which has $g_{\nu}\simeq 10^{-12}$. But such small value for $g_{\nu}$ implies a KK scale for the large dimension $m_{\rm KK}\simeq g_{\nu}^{3}M_{\rm P}\simeq 10^{-57}$ GeV, and a ridiculously small species scale $\Lambda\simeq m_{\rm KK}^{1/3}M_{\rm P}^{2/3}\simeq gM_{\rm P}\simeq 100$ eV. Thus a single dimension does not seem to work if you want to obtain Dirac neutrino masses from tiny Yukawa couplings.
A.2 One large dimension and the ‘Dark Dimension’ scenario
We have just seen that a single dimension does not work for the purpose of understanding the smallness of Dirac neutrino masses. Still we may asume that there exists a single large dimension because of other motivations. This is the case of the dark dimension scenario [57, 58, 59, 60]. In this scenario one imposes that there is a tower with $m_{\rm KK}\simeq V^{1/4}\simeq 10^{-12}$ GeV, related to the cosmological constant. Note that we saw that in this limit
$m_{\rm KK}\ \simeq\ \frac{M_{\rm P}}{u^{3/2}}\ ,$ (A.10)
so that in order to get $m_{\rm KK}\simeq 10^{-3}$ eV, one needs $u\simeq 10^{20}$ and $\Lambda\simeq 10^{8}$ GeV.
Let us now explore whether one can have a consistent structure for all the SM Yukawa couplings, although giving up on having a tiny Dirac neutrino mass. It is useful to study two cases separately depending on whether i) There is no tower of light gonions along the extra large dimension or ii) there is a tower of gonions. By gonion tower, we are referring here to intersections at which SM fields live.
i) One large dimension without light gonion tower
An example of this will be taking the limit $u^{(1)}\rightarrow\infty$ in the first complex plane in the toy model. It is easy to see that in this model in that limit the angles at the intersection of the matter fields $F_{R},F_{L},H$ remain finite, and hence there is no tower of light gonions (there are gonions at the $(aa^{*})$ intersections but the corresponding zero modes do not have Yukawa coupings to the rest). Let us consider then the Yukawa coupling $(F_{R}F_{L}H)$. It will have a structure
$Y_{{\cal N}=1}\ \simeq\ t^{1/4}e^{\phi_{4}/2}\left(\frac{m_{{\rm gon},1}}{M_{s}}\right)^{1/2}\left(\frac{m_{{\rm gon},2}}{M_{s}}\right)^{1/2}\left(\frac{m_{{\rm gon},3}}{M_{s}}\right)^{1/2}\ .$ (A.11)
Now, all gonions in all three planes will have masses of order $M_{s}$, so that the Yukawa couplings will be of order $Y\simeq t^{1/4}e^{\phi_{4}/2}\simeq 1$. Going to a more realistic SM setting as described in the main text, this is ok phenomenologically for the heaviest quark-lepton generation. Smaller Yukawa couplins for the first generations could come both from $(m_{\rm gon}/M_{s})^{1/2}$ powers or perhaps the holomorphic superpotential. [^6] It must be said that there is a point in which the structure of the toy model departures from a realistic SM configuration. Indeed, there are gonions at the $(aa^{*})$ intersections in the toy model. This fact comes along with a contribution to the $U(4)$ gauge coupling which decreases like $\sim 1/u^{1/2}$. This would not happen in a SM configuration in which $(aa^{*})$ intersections are absent, so there is no need for any SM gauge coupling to become too small.
So as a summary: a single large dimension along a direction which does not feature light gonions associated to the SM is compatible with the existence of a realistic structure of Yukawa couplings.
ii) One large dimension with light gonion towers
Consider now the case in which the large dimension goes along a complex plane in which a light tower of gonions develop. This is for example the case of the second plane in the toy model with $s,t^{(2)},u^{(2)}\sim u$, $u\rightarrow\infty$. The Yukawa couplings with a ‘${\cal N}=1$ structure’ would look like in eq.(A.11) above. However now the Yukawa is further supressed by a factor $({m_{{\rm gon},2}}/{M_{s}})^{1/2}\simeq 1/u^{1/4}\simeq 10^{-5}$. There are now two subcases, depending on whether the tower of gonions corresponds to right-handed neutrinos or to the rest of the SM particles:
1) The gonion towers correspond to right-handed neutrinos . In this case there will be a tower of $\nu_{R}$-like gonions at a scale $m_{gon,\nu}\simeq 10^{-10}M_{s}\simeq 10$ MeV. Neutrinos will get a Yukawa of order $Y_{\nu}\simeq 10^{-5}$. One does not expect a large instanton generated Majorana mass for the right-handed neutrinos, since they would be exponentially suppressed like $e^{-t}$. Thus a usual see-saw would not be available and perhaps some other mechanism involving the massive gonions could be at work. Asuming this is possible, indeed a large single dimension along a direction involving $\nu_{R}$ gonions could be consistent.
2) The gonion towers correspond to gonions of SM particles
In this case the above Yukawa would be too small to be consistent with the heaviest quark/lepton generations. So such type of Yukawa couplings would not be enough to describe the observed SM spectrum. Still, one can consider Yukawa couplings with a ‘${\cal N}=2$’ type of structure. In this case the Yukawa couplings will have the form
$Y_{{\cal N}=2}\,\simeq\,t^{1/4}g_{\rm SM}\left(\frac{m_{{\rm gon},1}}{M_{s}}\right)^{1/2}\left(\frac{m_{{\rm gon},3}}{M_{s}}\right)^{1/2}\,.$ (A.12)
Here the factor $g_{\rm SM}$ stands for a generic SM gauge coupling and comes from the prefactor $h_{22}^{-1/2}$ in eq.(3.7) in ref.[17]. Now Yukawa couplings as large as $Y\sim 1$, as required by the top quark, may be easily obtained. In fact the problem is the contrary since there is the prefactor $t^{1/4}\simeq u^{1/4}\sim 10^{5}$ which makes the Yukawa too large unless it is compensated by the gonion factors in the right. So, assuming SUSY, Yukawa couplings as large as one, but not larger, are possible if